# Nonparametric Least squares estimators for interval censoring

### v4

@@ 4 Conclusion
We stated a limit result for the least squares estimator minimizing (1.6) for interval censoring, case 2, in the “non-separated case” (“observation intervals can be arbitrariy small”), in which situation the limit behavior of the nonparametric MLE is still unknown. This is Theorem 1 in section 2. For the separated case the limit distribution of the nonparametric MLE was derived in [piet:96]. One could say (as has been done) that the separated case is the more important case (“there will always be a positive time interval between examination times of the doctor”), but the situation where one does not have the limit for the non-separated case is rather unsatisfactory. After all, allowing arbitrarily small intervals is mostly a matter of type of scaling one allows.
There is a conjecture that the nonparametric MLE has a faster rate of convergence in section 5.2 of part 2 of [GrWe:92] and there also has been the construction of a histogram-type estimator with a faster local rate in [lucien:99], but the result for the least squares estimator and simulations suggest that this conjecture may not hold for the MLE. See, in particular, Figure 5 for sample size $n=10,000$ in the present paper. But it is conceivable that the $(n\log n)^{1/3}$-rate starts showing up for astronomical sample sizes. It is true that for sample size $n=10,000$ its variance has become smaller than that of the LS estimate, reversing the relation between these variances for $n=1000$.
@@
C++ routines with makefiles to check the algorithms and simulations are given in [piet_intcens:25]. They use the iterative convex minorant algorithm.
@@ 5.1 Variance and drift of the relevant limit process for Theorem 1
Then we have, for fixed $t_{0}\in(0,M)$, and $W_{m,\hat{F}_{n}}$ is defined by (2)
@@
$\qquad\qquad\qquad\qquad+\int_{v\in[t_{0},t_{0}+n^{-1/3}t]}\psi_{2,F_{0}}(u,v,\delta_{0},\delta_{1})\,d\bigl({\mathbb{Q}}_{n}-Q_{0}\bigr),$
@@
Let the conditions of Theorem 1 be satisfied and let $\hat{F}_{n}$ be the least squares estimate, minimizing (2.8). Then $n^{2/3}X_{n}$ converges in distribution, in the Skorohod topology, to the process
$t\mapsto\sigma_{t_{0}}W(t),\qquad t\in{\mathbb{R}},$
where $W$ is standard two-sided Brownian motion and $\sigma_{t_{0}}$ is defined by (2.7).
@@
Note the similarity with Lemma 9.4 in [piet:24], but note that we now take $F_{0}$ instead of $\hat{F}_{n}$ in the definition of $X_{n}$.
@@
Note that we can also write:
@@
$n^{2/3}X_{n}\stackrel{{\scriptstyle{\cal D}}}{{\longrightarrow}}\sigma_{t_{0}}W,\qquad n\to\infty.$
@@
We can now write
$W_{n,\hat{F}_{n}}(t_{0}+n^{-1/3}t)-W_{n,\hat{F}_{n}}(t_{0})=X_{n}(t)+Y_{n}(t),$
where $Y_{n}$ is defined by
$Y_{n}(t)$ $=\int_{u\in[t_{0},t_{0}+n^{-1/3}t]}\left\{F_{0}(u)-\hat{F}_{n}(u)\right\}\,d{\mathbb{H}}_{n}(u,v)$
$\qquad-\int_{u\in[t_{0},t_{0}+n^{-1/3}t]}\left\{F_{0}(v)-F_{0}(u)-\hat{F}_{n}(v)+\hat{F}_{n}(u)\right\}\,d{\mathbb{H}}_{n}(u,v)$
$\qquad+\int_{v\in[t_{0},t_{0}+n^{-1/3}t]}\left\{F_{0}(v)-F_{0}(u)-\hat{F}_{n}(v)+\hat{F}_{n}(u)\right\}\,d{\mathbb{H}}_{n}(u,v)$
$\qquad+\int_{v\in[t_{0},t_{0}+n^{-1/3}t]}\left\{F_{0}(v)-\hat{F}_{n}(v)\right\}\,d{\mathbb{H}}_{n}(u,v),$ (5.6)
and where ${\mathbb{H}}_{n}$ is the empirical distribution function of the $(U_{i},V_{i})$.
Let the process $G_{n}$ be defined by
$G_{n}(t)=\int_{u\in[0,t]}\,d{\mathbb{H}}_{n}(u,v)+\int_{v\in[0,t]}\,d{\mathbb{H}}_{n}(u,v),\qquad t\geq 0.$
Then $\hat{F}_{n}$ is the left-continuous slope of the (“self-induced”) cusum diagram, consisting of the points $(0,0)$ and
$\left(G_{n}(t),\int_{w\in[0,t]}\hat{F}_{n}(w)\,dG_{n}(w)+W_{n,\hat{F}_{n}}(t)\right),\qquad t\geq 0,$
truncated at 0 and $1$ (if the slope at $t$ is <0<0, we put $\hat{F}_{n}(t)=0$ and similarly, if the slope at $t$ is >1>1, we put $\hat{F}_{n}(t)=1$).
We have:
$\int_{u\in[t_{0},t_{0}+n^{-1/3}t]}\hat{F}_{n}(w)\,dG_{n}(w)+Y_{n}(t)$
$=2\int_{u\in[t_{0},t_{0}+n^{-1/3}t]}F_{0}(u)\,d{\mathbb{H}}_{n}(u,v)+2\int_{v\in[t_{0},t_{0}+n^{-1/3}t]}F_{0}(u)\,d{\mathbb{H}}_{n}(u,v)$
@@
$\qquad-\int_{v\in[t_{0},t_{0}+n^{-1/3}t]}\left\{F_{0}(u)-\hat{F}_{n}(u)\right\}\,d{\mathbb{H}}_{n}(u,v).$ (5.7)
The last two terms are smooth functionals of the model and can be shown to be of order $O_{p}(n^{-5/6})$. We shall discuss this matter below in sections 5.2 and Section 3.
The conclusion is that
$\int_{u\in[t_{0},t_{0}+n^{-1/3}t]}\hat{F}_{n}(w)\,dG_{n}(w)+W_{n,\hat{F}_{n}}(t_{0}+n^{-1/3}t)-W_{n,\hat{F}_{n}}(t_{0})$
$=2\int_{u\in[t_{0},t_{0}+n^{-1/3}t]}F_{0}(u)\,d{\mathbb{H}}_{n}(u,v)+2\int_{v\in[t_{0},t_{0}+n^{-1/3}t]}F_{0}(v)\,d{\mathbb{H}}_{n}(u,v)$
$\qquad+X_{n}(t)+O_{p}(n^{-5/6}).$
This means that $n^{1/3}\{\hat{F}_{n}(t_{0})-F_{0}(t_{0})\}$ behaves asymptotically as the slope of the greatest convex minorant of the process
$2n^{2/3}\int_{u\in[t_{0},t_{0}+n^{-1/3}t]}\{F_{0}(u)-F_{0}(t_{0})\}\,d{\mathbb{H}}_{n}(u,v)$
$\qquad+2n^{2/3}\int_{v\in[t_{0},t_{0}+n^{-1/3}t]}\{F_{0}(u)-F_{0}(t_{0})\}\,d{\mathbb{H}}_{n}(u,v)+n^{2/3}X_{n}(t)+O_{p}(n^{-1/6}).$
in the time scale
$n^{1/3}\{G_{n}(t_{0}+n^{-1/3})-G_{n}(t_{0})\}\sim t\{h_{1}(t)+h_{2}(t)\}.$
@@
$2n^{2/3}\int_{u\in[t_{0},t_{0}+n^{-1/3}t]}\{F_{0}(u)-F_{0}(t_{0})\}\,d{\mathbb{H}}_{n}(u,v)$
$\qquad+2n^{2/3}\int_{v\in[t_{0},t_{0}+n^{-1/3}t]}\{F_{0}(u)-F_{0}(t_{0})\}\,d{\mathbb{H}}_{n}(u,v)$
$\sim f_{0}(t_{0})\{h_{1}(t_{0})+h_{2}(t_{0})\}t^{2},$
the result now follows from Brownian scaling. using a similar computation on the interval $t_{0}+n^{-1/3}t,t_{0}]$ for $t<0$.
@@ 5.2 Negligibility of the “off-diagonal terms”
In (5.1) we met two terms which we call “off-diagonal terms”, because, for example in the interior point method for solving the minimization problem, they correspond to non-zero off-diagonal terms of the Hessian. This is a difficulty we do not have to deal with in the current status model. They are the terms

### v5

@@ 4 Conclusion
We stated a limit result for the least squares estimator minimizing (1.6) for interval censoring, case 2, in the “non-separated case” (“observation intervals can be arbitrariy small”), in which situation the limit behavior of the nonparametric MLE is still unknown. This is Theorem 1 in section 2. For the separated case the limit distribution of the nonparametric MLE was derived in [6]. One could say (as has been done) that the separated case is the more important case (“there will always be a positive time interval between examination times of the doctor”), but the situation where one does not have the limit for the non-separated case is rather unsatisfactory. After all, allowing arbitrarily small intervals is mostly a matter of type of scaling one allows.
There is a conjecture that the nonparametric MLE has a faster rate of convergence in section 5.2 of part 2 of [4] and there also has been the construction of a histogram-type estimator with a faster local rate in [2], but this conjecture does not show up for the sample sizes $n=1000$ and $n=10,000$ in our simulations. See, in particular, Figure 5 for sample size $n=10,000$ in the present paper. But it is conceivable that the $(n\log n)^{1/3}$-rate starts showing up for astronomical sample sizes. It is true that for sample size $n=10,000$ its variance has become smaller than that of the LS estimate, reversing the relation between these variances for $n=1000$.
@@
C++ routines with makefiles to check the algorithms and simulations are given in [9]. They use the iterative convex minorant algorithm.
@@ 5.1 Variance and drift of the relevant limit process for Theorem 1
Then we have, for fixed $t_{0}\in(0,M)$, and $W_{n,\hat{F}_{n}}$ is defined by (2)
@@
$\qquad\qquad\qquad\qquad+\int_{v\in[t_{0},t_{0}+n^{-1/3}t]}\psi_{2,F_{0}}(u,v,\delta_{0},\delta_{1})\,d\bigl({\mathbb{Q}}_{n}-Q_{0}\bigr),$ (5.4)
@@
Let the conditions of Theorem 1 be satisfied. Then $n^{2/3}X_{n}$ converges in distribution, in the topology of uniform convergence on compacta, to the process
$t\mapsto a_{t_{0}}W(t),\qquad t\in{\mathbb{R}},$
where $W$ is standard two-sided Brownian motion and $a_{t_{0}}$ is defined by (2).
@@
Note the similarity with Lemma 9.4 in [8], but note that we now take $F_{0}$ instead of $\hat{F}_{n}$ in the definition of $X_{n}$.
@@
Note that we can write (5.1) in the form:
@@
$n^{2/3}X_{n}\stackrel{{\scriptstyle{\cal D}}}{{\longrightarrow}}a_{t_{0}}W,\qquad n\to\infty.$
@@
5.2 Proof of Theorem 1
Proof.
We now generalize the approach in the proof of Theorem 3.7 in [7]. Let ${\mathbb{G}}_{n}$ be defined by
${\mathbb{G}}_{n}(t)=2\int_{u\leq t}\,d{\mathbb{H}}_{n}(u,v)+2\int_{v\leq t}\,d{\mathbb{H}}_{n}(u,v),$
and let $V_{n}$ be defined by
$V_{n}(t)=\int_{w\leq t}\hat{F}_{n}(w)\,d{\mathbb{G}}_{n}(w)+W_{n,\hat{F}_{n}}(t),\qquad t\geq 0,$
where $W_{n,\hat{F}_{n}}$ is defined by (2). The values of the processes are zero for $t\leq 0$.
Then $\hat{F}_{n}$ is the left-continuous slope of the greatest convex minorant of cumulative sum (“cusum”) diagram
$({\mathbb{G}}_{n}(t),V_{n}(t)),\qquad t\geq 0,$ (5.7)
where (5.7) runs through the points $({\mathbb{G}}_{n}(U_{i}),V_{n}(U_{i}))$ and ${\mathbb{G}}_{n}(V_{i}),V_{n}(V_{i}))$, $i=1,\dots,n$, and the slopes are put equal to zero if they are less than zero and equal to $1$ if they are larger than $1$.
We next define, for $a\in(0,1)$,
$U_{n}(a)=\text{argmin}\left\{t\in{\mathbb{R}}:V_{n}(t)-a{\mathbb{G}}_{n}(t)\right\},$
where we take the supremum in case of multiple argmins and use, for $a_{0}=F_{0}(t_{0})$, the “switch relation”
$P\left\{n^{1/3}\bigl\{\hat{F}_{n}(t_{0})-F_{0}(t_{0})\bigr\}\geq x\right\}=P\left\{U_{n}(a_{0}+n^{-1/3}x)\leq t_{0}\right\},$
see Figure 3.7 on p. 69 of [7]. We have:
$U_{n}(a_{0}+n^{-1/3}x)-t_{0}=\text{argmin}\left\{t\in{\mathbb{R}}:V_{n}(t)-(a_{0}+n^{-1/3}x){\mathbb{G}}_{n}(t)\right\}-t_{0}$
$=\text{argmin}\left\{t_{0}+n^{-1/3}t\in{\mathbb{R}}:V_{n}(t_{0}+n^{-1/3}t)-(a_{0}+n^{-1/3}x){\mathbb{G}}_{n}(t_{0}+n^{-1/3}t)\right\}-t_{0}$
$=n^{-1/3}\text{argmin}\left\{t\in{\mathbb{R}}:V_{n}(t_{0}+n^{-1/3}t)-(a_{0}+n^{-1/3}x){\mathbb{G}}_{n}(t_{0}+n^{-1/3}t)\right\}.$
So we can conclude
$n^{1/3}\bigl\{U_{n}(a_{0}+n^{-1/3}x)-t_{0}\bigr\}$
$=\text{argmin}\left\{t\in{\mathbb{R}}:V_{n}(t_{0}+n^{-1/3}t)-(a_{0}+n^{-1/3}x){\mathbb{G}}_{n}(t_{0}+n^{-1/3}t)\right\},$
and we have to determine
$P\left\{n^{1/3}\bigl\{U_{n}(a_{0}+n^{-1/3}x)-t_{0}\bigr\}\leq 0\right\}.$
Since the argmin function does not change by adding a constant to the argument or multiplying the argument by a constant, we can replace $n^{1/3}\left\{U_{n}(a_{0}+n^{-1/3}x)-t_{0}\right\}$ by the argmin (in $t$) of:
$n^{2/3}\bigl\{V_{n}(t_{0}+n^{-1/3}t)-V_{n}(t_{0})\bigr\}-n^{2/3}(a_{0}+n^{-1/3}x)\bigl\{{\mathbb{G}}_{n}(t_{0}+n^{-1/3}t)-{\mathbb{G}}_{n}(t_{0})\bigr\}.$
We get, if $t\geq 0$,
$V_{n}(t_{0}+n^{-1/3}t)-V_{n}(t_{0})-a_{0}\left\{{\mathbb{G}}_{n}(t_{0}+n^{-1/3}t)-{\mathbb{G}}_{n}(t_{0})\right\}$
$=\int_{w\in(t_{0},t_{0}+n^{-1/3}t]}\left\{\hat{F}_{n}(w)-F_{0}(t_{0})\right\}\,d{\mathbb{G}}_{n}(w)+W_{n,\hat{F}_{n}}(t_{0}+n^{-1/3}t)-W_{n,\hat{F}_{n}}(t_{0})$
$=\int_{w\in(t_{0},t_{0}+n^{-1/3}t]}\left\{F_{0}(w)-F_{0}(t_{0})\right\}\,d{\mathbb{G}}_{n}(w)+X_{n}(t)$
@@ 5.2 Proof of Theorem 1
$\qquad-\int_{v\in[t_{0},t_{0}+n^{-1/3}t]}\left\{F_{0}(u)-\hat{F}_{n}(u)\right\}\,d{\mathbb{H}}_{n}(u,v),$
where $X_{n}$ is defined by (5.1). The last two terms are smooth functionals of the model and can be shown to be of order $O_{p}(n^{-5/6})$. We discuss this matter in section 5.3 and Section 3.
So we get:
$n^{2/3}\left\{V_{n}(t_{0}+n^{-1/3}t)-V_{n}(t_{0})\right\}-n^{2/3}(a_{0}+n^{-1/3}x)\left\{{\mathbb{G}}_{n}(t_{0}+n^{-1/3}t)-{\mathbb{G}}_{n}(t_{0})\right\}$
$=n^{2/3}\int_{w\in(t_{0},t_{0}+n^{-1/3}t]}\left\{F_{0}(w)-F_{0}(t_{0})\right\}\,d{\mathbb{G}}_{n}(w)+n^{2/3}X_{n}(t)$
$\qquad\qquad\qquad\qquad-n^{1/3}x\left\{{\mathbb{G}}_{n}(t_{0}+n^{-1/3}t)-{\mathbb{G}}_{n}(t_{0})\right\}+O_{p}\left(n^{-1/6}\right).$ (5.8)
@@
$n^{2/3}\int_{u\in[t_{0},t_{0}+n^{-1/3}t]}\{F_{0}(w)-F_{0}(t_{0})\}\,d{\mathbb{G}}_{n}(w)\sim f_{0}(t_{0})\{h_{1}(t_{0})+h_{2}(t_{0})\}t^{2},$
the right-hand side of (5.2) converges in distribution, in the topology of uniform convergence on compacta, to the process
$t\mapsto a_{t_{0}}W(t)+f_{0}(t_{0})\{h_{1}(t_{0})+h_{2}(t_{0})\}t^{2}-2\{h_{1}(t_{0})+h_{2}(t_{0})\}x,\qquad t\geq 0.$
We get a similar result for $t\leq 0$, we and get therefore:
$P\left[n^{1/3}\left\{U_{n}(a_{0}+n^{-1/3}x)-t_{0}\right\}\leq 0\right]$
$\longrightarrow P\left[\text{argmin}_{t}\left\{a_{t_{0}}W(t)+f_{0}(t_{0})\{h_{1}(t_{0})+h_{2}(t_{0})\}t^{2}-2\{h_{1}(t_{0})+h_{2}(t_{0})\}x\right\}\leq 0\right].$
Next note that
$P\left[\text{argmin}_{t}\left\{a_{t_{0}}W(t)+f_{0}(t_{0})\{h_{1}(t_{0})+h_{2}(t_{0})\}t^{2}-2\{h_{1}(t_{0})+h_{2}(t_{0})\}x\right\}\leq 0\right]$
$=P\left[\text{argmin}_{t}\left\{a_{t_{0}}W(t)+f_{0}(t_{0})\{h_{1}(t_{0})+h_{2}(t_{0})\}\left(t-\frac{x}{f_{0}(t_{0})}\right)^{2}\right\}\leq 0\right]$
Defining
$T(x)=\text{argmin}_{t}\left\{a_{t_{0}}W(t)+f_{0}(t_{0})\{h_{1}(t_{0})+h_{2}(t_{0})\}\left(t-\frac{x}{f_{0}(t_{0})}\right)^{2}\right\}-\frac{x}{f_{0}(t_{0})}\,$,
we have that the process
$x\mapsto T(x),\qquad x\in{\mathbb{R}},$
is stationary (this is one of the key results of [5]), so we have:
$P\left[\text{argmin}_{t}\left\{a_{t_{0}}W(t)+f_{0}(t_{0})\{h_{1}(t_{0})+h_{2}(t_{0})\}\left(t-\frac{x}{f_{0}(t_{0})}\right)^{2}\right\}\leq 0\right]$
$=P\left\{T(x)\leq 0\right\}=P\left\{T(0)\leq-\frac{x}{f_{0}(t_{0})}\right\}=P\left\{T(0)\geq\frac{x}{f_{0}(t_{0})}\right\},$
where the last equality follows from. the symmetry of the distribution of $T(0)$ around zero (compare with Theorem 5.2, p. 95 in part II of [4]).
The result of Theorem 1 now follows from Brownian scaling, ∎
@@ 5.3 Negligibility of the “off-diagonal terms”
In subsection 5.2 we met two terms which we call “off-diagonal terms”, because, for example in the interior point method for solving the minimization problem, they correspond to non-zero off-diagonal terms of the Hessian. This is a difficulty we do not have to deal with in the current status model. They are the terms
@@
5.4 Theorem 2
The proof proceeds along the lines of the proof of Theorem 1, but is much simpler, because we do not have to deal with “off-diagonal terms” and can compute the estimator in one step. Iterations, as in the situation of Theorem 1, are not needed for the computation and the cusum diagram is not “self-induced”. As noted above, the price we have to pay for this simplicity, though, is that the estimator is less efficient.
We start by analyzing the covariance structure of the process $W_{n,F_{0}}^{(2)}$, defined by (2). Analogously to the proof of Theorem 1, we define the process $X_{n}^{(2)}$ by
$X_{n}^{(2)}(t)=W_{n,F_{0}}^{(2)}(t_{0}+n^{-1/3}t)-W_{n,F_{0}}^{(2)}(t_{0}).$
For the process $X_{n}^{(2)}$ we have the following lemma.
Lemma 6.
Let the conditions of Theorem 1 be satisfied. Then $n^{2/3}X_{n}^{(2)}$ converges in distribution, in the topology of uniform convergence on compacta, to the process
$t\mapsto a_{t_{0}}^{\prime}W(t),\qquad t\in{\mathbb{R}},$
where $W$ is standard two-sided Brownian motion and $a_{t_{0}}^{\prime}$ is defined by (2.10).
Proof.
Let $t\geq 0$. Then:
$X_{n}^{(2)}(t)$ $=n^{-1}\sum_{j:t_{0}\leq U_{j}\leq t_{0}+n^{-1/3}t}\left\{\Delta_{j0}-F_{0}(U_{j})\right\}$
$\qquad\qquad\qquad\qquad+n^{-1}\sum_{j:t_{0}\leq V_{j}\leq t_{0}+n^{-1/3}t}\left\{\Delta_{j0}+\Delta_{j1}-\{F_{0}(V_{j})\}\right\}.$ (5.11)
This means:
$n^{4/3}\text{var}\left(X_{n}^{(2)}(t)\right)\sim F_{0}(t_{0})\{1-F_{0}(t_{0})\}\{h_{1}(t_{0})+h_{2}(t_{0})\}.$
The covariance between $\Delta_{j0}$ and $\Delta_{j0}+\Delta_{j1}$ in the first and second line of (5.4) gives a contribution of lower order, since this involves a $U_{j}$ and $V_{j}$ both lying in the shrinking interval $t_{0},t_{0}+n^{-1/3}t]$, which has a probability tending to zero, as $n\to\infty$.
So the result follows again from tightness and the central limit theorem. ∎
Proof of Theorem 2.
This time we define
${\mathbb{G}}_{n}(t)=\int_{u\leq t}\,d{\mathbb{H}}_{n}(u,v)+\int_{v\leq t}\,d{\mathbb{H}}_{n}(u,v),$
and, as in the proof of Theorem 1, $V_{n}$ by
$V_{n}(t)=\int_{w\leq t}\hat{F}_{n}(w)\,d{\mathbb{G}}_{n}(w)+W_{n,\hat{F}_{n}}^{(2)}(t),\qquad t\geq 0.$
The values of the processes are zero for $t\leq 0$. Note that $V_{n}$ can this time be written in the form
$V_{n}(t)=\int_{w\leq t}F_{0}(w)\,d{\mathbb{G}}_{n}(w)+W_{n,F_{0}}^{(2)}(t),\qquad t\geq 0,$
so $\hat{F}_{n}$ is not part of this representation of $V_{n}$.
We next define, for $a\in(0,1)$,
$U_{n}(a)=\text{argmin}\left\{t\in{\mathbb{R}}:V_{n}(t)-a{\mathbb{G}}_{n}(t)\right\},$
taking the supremum in case of multiple argmins and use, for $a_{0}=F_{0}(t_{0})$, the switch relation
$P\left\{n^{1/3}\bigl\{\hat{F}_{n}(t_{0})-F_{0}(t_{0})\bigr\}\geq x\right\}=P\left\{U_{n}(a_{0}+n^{-1/3}x)\leq t_{0}\right\}.$
We have again
$n^{1/3}\left\{U_{n}(a_{0}+n^{-1/3}x)-t_{0}\right\}$
$=\text{argmin}\left\{t\in{\mathbb{R}}:V_{n}(t_{0}+n^{-1/3}t)-(a_{0}+n^{-1/3}x){\mathbb{G}}_{n}(t_{0}+n^{-1/3}t)\right\},$
and we have to determine
$P\left\{n^{1/3}\left\{U_{n}(a_{0}+n^{-1/3}x)-t_{0}\right\}\leq 0\right\}.$
We get, if $t\geq 0$,
$V_{n}(t_{0}+n^{-1/3}t)-V_{n}(t_{0})-a_{0}\left\{{\mathbb{G}}_{n}(t_{0}+n^{-1/3}t)-{\mathbb{G}}_{n}(t_{0})\right\}$
$=\int_{w\in(t_{0},t_{0}+n^{-1/3}t]}\left\{F_{0}(w)-F_{0}(t_{0})\right\}\,d{\mathbb{G}}_{n}(w)+X_{n}^{(2)}(t).$
Hence:
$n^{2/3}\left\{V_{n}(t_{0}+n^{-1/3}t)-V_{n}(t_{0})\right\}-n^{2/3}(a_{0}+n^{-1/3}x)\left\{{\mathbb{G}}_{n}(t_{0}+n^{-1/3}t)-{\mathbb{G}}_{n}(t_{0})\right\}$
$=n^{2/3}\int_{w\in(t_{0},t_{0}+n^{-1/3}t]}\left\{F_{0}(w)-F_{0}(t_{0})\right\}\,d{\mathbb{G}}_{n}(w)+n^{2/3}X_{n}^{(2)}(t)$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad-n^{1/3}x\left\{{\mathbb{G}}_{n}(t_{0}+n^{-1/3}t)-{\mathbb{G}}_{n}(t_{0})\right\}.$ (5.12)
Since
$n^{2/3}\int_{u\in[t_{0},t_{0}+n^{-1/3}t]}\{F_{0}(w)-F_{0}(t_{0})\}\,d{\mathbb{G}}_{n}(w)\sim\tfrac{1}{2}f_{0}(t_{0})\{h_{1}(t_{0})+h_{2}(t_{0})\}t^{2},$
the right-hand side of (5.4) converges in distribution, in the topology of uniform convergence on compacta, to the process
$t\mapsto a_{t_{0}}^{\prime}W(t)+\tfrac{1}{2}f_{0}(t_{0})\{h_{1}(t_{0})+h_{2}(t_{0})\}t^{2}-\{h_{1}(t_{0})+h_{2}(t_{0})\}x,\qquad t\geq 0.$
We get a similar result for $t\leq 0$, we and get therefore:
$P\left[n^{1/3}\left\{U_{n}(a_{0}+n^{-1/3}x)-t_{0}\right\}\leq 0\right]$
$\longrightarrow P\left[\text{argmin}_{t}\left\{a_{t_{0}}^{\prime}W(t)+\tfrac{1}{2}f_{0}(t_{0})\{h_{1}(t_{0})+h_{2}(t_{0})\}t^{2}-\{h_{1}(t_{0})+h_{2}(t_{0})\}x\right\}\leq 0\right].$
Next note that
$P\left[\text{argmin}_{t}\left\{a_{t_{0}}^{\prime}W(t)+f_{0}(t_{0})\{h_{1}(t_{0})+h_{2}(t_{0})\}t^{2}-\{h_{1}(t_{0})+h_{2}(t_{0})\}x\right\}\leq 0\right]$
$=P\left[\text{argmin}_{t}\left\{a_{t_{0}}^{\prime}W(t)+\tfrac{1}{2}f_{0}(t_{0})\{h_{1}(t_{0})+h_{2}(t_{0})\}\left(t-\frac{x}{f_{0}(t_{0})}\right)^{2}\right\}\leq 0\right]$
From this point on the proof can be completed in the same way as the proof of Theorem 1 with as only differences the presence of the factor $\tfrac{1}{2}$ in front of $f_{0}(t_{0})$ and the scale factor $a_{t_{0}}^{\prime}$ instead of $a_{t_{0}}$. ∎